### v1

@@ 3.3 Documentation of functions
The purpose of this section is to provide a more comprehensive documentation of all the most important functions included in [Uncaptioned image] and their options. All these functions are internally memorized to improve performance, and in Sec. 4 we provide concrete examples that illustrate their usage.
@@ 3.3.1 SOFIA command
[…]
@@
[…]
@@
Subtopology history graph.
When IncludeSubtopologies -> True, the option ShowHistoryGraph can be used to print a top-bottom graph (laid out based on its spanning tree) whose nodes represent (sub)topologies, with edges connecting those related by edge contractions. If ShowHistoryGraph -> True, both the diagrams and the corresponding singularities are displayed at each node (see Figs. 2(a) and 2(b), respectively); if ShowHistoryGraph -> SingOnly, only the singularities are shown.
Figure 2: Example output of SOFIA[{ edges, nodes} , ShowHistoryGraph -> True, SolverBound -> Infinity] applied to the parachute diagram with edges = { { { 1,2} ,m} , { { 1,2} ,m} , { { 2,3} ,m} , { { 1,3} ,m} } and nodes = { { $\mathtt{M_{1}}$} , { $\mathtt{M_{2}}$} , { $\mathtt{M_{3}}$} } . Panel (a) illustrates how diagrams are connected through edge contractions (with tadpoles removed), while panel (b) shows the corresponding candidate singularities. As explained in the main text, SOFIA[] uses the shorthands $\mathtt{mm=m^{2}}$ and $\mathtt{MMi=M_{i}^{2}}$ in its outputs. Note that the singularities in (b) are in one-to-one correspondence with those in the database [17]—that is, there are no fictitious singularities in this example.
@@
[…]
@@
[…]
Lines 13–21 give options that control the solving strategy.
@@
Within [Uncaptioned image] , there are two available solvers: FastFubini (which implements the analytic procedure outlined in Sec. 2.2) and momentumPLD (which instead uses the numerical approach of Sec. 2.2). The former relies solely on Mathematica functions, while the latter requires the working Julia dependencies detailed in App. B. All the options PLDname relate specifically to momentumPLD (lines 13–18) are described under the same name as in [65]. Finally, the option FactorResult -> False can be used to skip the factorization into irreducible components of $\mathcal{S}(\vec{s})$ in (2.24), which, for large examples, can require both a long runtime and a large amount of memory.
[…]
Lines 22–24 set options for Effortless, which are relevant only when FindLetters -> True is enabled.
@@
[…]
@@ 3.3.4 Graphical functions
[…]
@@
[…]
@@
For example, Fig. 3 displays two pairs of diagrams (a-b and c-d) related by symmetry. Running Symmetry[(a),(b)]—with (a) and (b) representing the lists of edges and nodes of the diagrams of the same names—gives { (a),{ $\mathtt{s_{1,2}\!\to\!s_{2,3},s_{2,3}\!\to\!s_{1,2},M_{1}^{2}\!\to\!M_{2}^{2},M_{4}^{2}\!\to\!M_{1}^{2}}$} } , which corresponds to a clockwise rotation of $90^{\circ}$ relating diagram (a) to diagram (b).
Diagrams (c) and (d) have a less obvious relation. Running Symmetry[(c),(d)] outputs { (d),{ $\mathtt{s_{1,2}\!\to\!0,s_{2,3}\!\to\!0,M^{2}\to s_{1,2}}$} } which effectively amounts to setting the momentum $p_{4}$ of diagram (d) to zero, $p_{4}\to 0$. Indeed, in terms of Mandelstam variables, we have $\,p_{4}\cdot p_{1}=\,p_{2}\cdot p_{3}\propto s_{23}\to 0$ and $\,p_{4}\cdot p_{3}=\,p_{1}\cdot p_{2}\propto s_{12}\to 0$. In this case, Symmetry[] also correctly recognizes that the mass scale $M^{2}$ in diagram (d) plays the role of the Mandelstam invariant $s_{34}=s_{12}$ in diagram (c).
Figure 3: Diagrams related by the Symmetry function as described in the text. These diagrams were all generated by the FeynmanPlot function. The dashed lines denote massless particles and otherwise the mass label is displayed on the edges.
@@
In such cases, kinematic replacements can be subtle because the operations of specializing kinematics and computing singularities generally do not commute (a phenomenon analogous to that encountered when specializing kinematics and evaluating Feynman integrals; see [66, 67, 15]). To address this, we introduce a small parameter $\delta$ and extract the leading non-vanishing coefficient for each component of the singular locus $\mathcal{S}$ as $\delta\to 0$. See App. C for an explicit demonstration.
[…]
@@ 4 Detailed examples
We now demonstrate the explicit use of the functions introduced above with concrete examples that are not too large to fit within the text. Collectively, these examples took less than 30 seconds to run on a conventional laptop. A list of 30 additional examples, including many multi-scale and multi-loop ones, can be found in the Mathematica notebook SOFIA_ examples.nb [3].
@@ 4.1 Symbol alphabet
In this example, we illustrate how [Uncaptioned image] , combined with the publicly available package Effortless [36], can be used to determine a (super)set of logarithmic letters, both even and odd, for a given Feynman integral.
We consider the diagram in Fig. 4(a), which contributes to the $t$-channel planar Møller and Bhabha scattering amplitudes at $\mathcal{O}(e^{6})$ in QED. The massive electrons and/or positrons are represented by solid lines, while the massless photon exchanges are depicted by dashed lines. This diagram was computed explicitly using differential equations in [68] and thus serves as an excellent crosscheck to the [Uncaptioned image] implementation.
@@
The output is a (super)set of 10 even and 42 odd candidate letters, which we collect in the alphabet SOFIAalphabet[[1;;2]]=$=\{\{\tilde{W}_{1},\ldots,\tilde{W}_{10}\},\{\tilde{W}_{11},\ldots,\tilde{W}_{52}\}\}$. We use the notation $\tilde{W}_{i}=\log(W_{i})$. For example:
$\tilde{W}_{31}$ = Log[$\frac{\mathtt{s12\,(4mm^{2}+mm\,s12-s12\,s23)+r[[5]]\,r[[10]]}}{\mathtt{s12\,(4mm^{2}+mm\,s12-s12\,s23)-r[[5]]\,r[[10]]}}$];
$\tilde{W}_{32}$ = Log[$\frac{\mathtt{s12\,(4mm-s12)(3mm-s23)+r[[5]]\,r[[10]]}}{\mathtt{s12\,(4mm-s12)(3mm-s23)-r[[5]]\,r[[10]]}}$];
@@ 2 Sqrt[4 mm^3 - mm^2 s12 + 2 mm s12 s23 - s12 s23^2],
As we shall show momentarily, $\mathtt{r}$ is sufficient for the SOFIA command to produce a complete basis of odd letters. (Note that the signs chosen in the DoubleRoots option above are purely conventional and ensure that the comparison below is performed using the root conventions of [68]).
At this stage, we can check the validity of SOFIAalphabet by comparing it against the alphabet knownAlphabet in [68, Eq. (2.8)]. This is done by running:
@@
In line 1 above, the syntax reflects that the roots defined in [68, Eq. (2.6)] were manually re-expressed, for comparison purposes, in terms of the roots $\mathtt{r[[i]]}$. The SOFIADecomposeAlphabet command then checks whether it is possible to decompose each element of knownAlphabet in terms of SOFIAalphabet and returns the decomposition if it is. For example, the most complicated letter in knownAlphabet decomposes according to:
$\mathrm{d}\log\frac{r_{5}r_{10}\left(m^{2}{-}s_{12}\right){-}4m^{6}s_{12}{-}3m^{4}s_{12}^{2}{+}m^{2}s_{12}^{3}{+}3m^{2}s_{12}^{2}s_{23}{-}s_{12}^{3}s_{23}}{r_{5}r_{10}\left(m^{2}{-}s_{12}\right){+}4m^{6}s_{12}{+}3m^{4}s_{12}^{2}{-}m^{2}s_{12}^{3}{-}3m^{2}s_{12}^{2}s_{23}{+}s_{12}^{3}s_{23}}=\mathrm{d}\tilde{W}_{31}{-}2\mathrm{d}\tilde{W}_{32}\,.$ (4.1)
See the Mathematica notebook SOFIA_ examples.nb [3] for all the details of the complete decomposition, as well as how the roots $r_{i}=\mathtt{r[[i]]}$ in r and the letters $\tilde{W}_{j}$ in SOFIAalphabet are defined. In the same notebook, we show that the same procedure successfully reproduces other non-trivial alphabets, including those arising in the two-loop massless and one-mass pentagon functions [69, 70, 71] as well as the genuine two-loop massless six-particle letters [72, 54].
@@ 4.2 Feynman geometries
An important early step in computing this amplitude is to determine the types of functions to which the contributing diagrams are expected to evaluate, i.e., the function space. This information is encoded in the underlying geometry of the associated master integrals. To reveal it, it suffices to carefully analyze the maximal cut of each master integral in four dimensions ($\varepsilon=0$), which is conveniently done in the Baikov representation thanks to (2.8). This approach is reliable because the underlying geometry is characterized by the solution of the homogeneous part of the differential equation satisfied by the uncut integrals, which coincides with the differential equation governing the maximal cut integral [73, 74].
The maximal cut procedure always reduces Feynman integrals to periods over an algebraic variety. The geometry of this variety dictates the class of functions that appear after integration. For instance, a genus-0 (rational) curve gives rise to polylogarithms, a genus-1 (elliptic) curve to elliptic functions, and more intricate cases may involve higher-genus Riemann surfaces (see, e.g., [75]) or even higher-dimensional Calabi-–Yau manifolds (see Sec.,4.3), pointing to increasingly rich—and still not fully understood—mathematical structures. Strictly speaking, this way, one can only place an upper bound on the complexity of a given Feynman geometry; verifying that it is not degenerate generally requires further analysis. (To our knowledge, there is currently no automated algorithm to systematically determine the precise function space of a Feynman integral, e.g., whether it is polylogarithmic, elliptic, or beyond. Even less is known at the amplitude level, where full cancellations among non-polylogarithmic sectors may occur; see, e.g., [76]).
@@
From (4.2), it is also easy to read off the leading singularities of the diagram in Fig. 4(b). Indeed, these are given by the zeroes of the discriminant of $P_{4}(x_{7})$ in $x_{7}$, i.e.,
@@
One can verify that all the singularities in the $\varepsilon$-factorized differential equation from [77] are included in the resulting list available in [3].
@@ 4.3 Eikonal/Regge/heavy-mass physics
This example considers the eikonal/Regge scattering of two heavy masses—e.g., black holes or neutron stars, treated as point particles—via graviton exchange. In particular, we focus on the diagram shown in Fig. 4(c), which contributes to the process at $\mathcal{O}(G_{N}^{5})$. This example demonstrates two key points: first, how to take the eikonal limit of the Baikov representation; and second, how the latter turns out sufficient to uncover the underlying Calabi–Yau threefold geometry [78, 42].
In order to take the eikonal/Regge/heavy-mass expansion and keep the leading order term which captures the relevant classical physics [79], we first parametrize the momenta in Fig. 4(c) such that
@@
The curve $\hat{P}_{4}(x_{12},x_{13},\hat{x}_{14})=0$ defines a Calabi–Yau threefold (as a quartic polynomial in $\mathbb{CP}^{3}$ [80]), in agreement with the results of [78, 42].
@@ 4.4.1 Energy-energy correlators
As an example, we focus on the type of integrals arising in the computation of the four-point energy-energy correlator (EEC) in the collinear limit within $\mathcal{N}=4$ super Yang-Mills theory [81]. Physically, the EEC captures meaningful correlations of energy flux after the emission of four final-state particles from an initial energetic parton and is now an active topic of research in both theoretical and experimental high-energy physics, see, e.g.[82, 83].
For four-point EEC, the relevant integrals to compute originate from a phase-space integral over a squared $1\to 4$ amplitude [81, Eqs. (2-4)]. For our discussion, it is sufficient to quote the two families of integrals considered in [81]:
@@
We treat $z$, $\bar{z}$, $w$, $\bar{w}$ as independent variables. The parametrization of these families is overcomplete since, e.g., $a_{2}$ and $a_{3}$ in (4.10) cannot be simultaneously positive. This would introduce an overlapping pole $\frac{1}{x_{12}x_{23}}$ which is forbidden by a version of Steinmann relations. We thus find it simpler to perform the singularity analysis at the level of master integrals introduced in [81, App. B] (in the same way we typically do the analysis at the level of individual Feynman diagram topologies, as opposed to all integrals at a given loop order at the same time).
The complete analysis is given in [3]. Below, we illustrate the procedure on a simple master integral
@@
[…]
Looping this command over all the master integrals in [81, App. B] immediately reproduces the polynomial part of the alphabet quoted in that reference, namely
@@
returns a distinctly complicated letter (see SOFIA_ examples.nb [3] for details):
@@
It has an interesting geometric interpretation. As explained in [81], the solutions to the maximal cut conditions of the integrals in (4.16) (i.e., $s_{123}=s_{234}=x_{1234}=0$) define a singular cubic curve after projection onto the $(x_{2},x_{3})$-plane. Its roots $a\,,b\,,c$, which give rise to the non-integer polynomial part of the alphabet listed in [81, Eq. (12)], satisfy

### v2

@@ 3.3 Documentation of functions
The purpose of this section is to provide a more comprehensive documentation of all the most important functions included in [Uncaptioned image] and their options. All these functions are memoized (that is, the results are stored and can be reused for future inputs, without being recomputed) to improve performance, and in Sec. 4 we provide concrete examples that illustrate their usage.
@@ 3.3.1 SOFIA command
LoopEdges specifies which internal edges are assigned $\ell_{1},\dots,\ell_{L}$ before [Uncaptioned image] searches for both an ordering and linear shifts in the $\ell$’s minimizing the number $N$ of Baikov integration variables. For example, LoopEdges -> {3,7} means $\ell_{1}$ and $\ell_{2}$ are initially assigned to the $3^{\text{rd}}$ and $7^{\text{th}}$ edge in the list edges. If LoopEdges -> Automatic, this choice is automatically made by the package to minimize $N$. V
@@
Negative powers of ISPs are included in (2.20) if IncludeISPs -> True and are excluded otherwise. In other words, IncludeISPs -> True places ISPs in the denominators, while IncludeISPs -> False places them in the numerators, where they do not contribute to singularities. V
@@
The options StartAtSubtopology -> n and EndAtSubtopology -> m allow the user to start and end the singularity analysis on the n-th and m-th subtopologies, respectively. V
Subtopology history graph.
When IncludeSubtopologies -> True, the option ShowHistoryGraph can be used to print a top-bottom graph (laid out based on its spanning tree) whose nodes represent (sub)topologies, with edges connecting those related by edge contractions. If ShowHistoryGraph -> True, both the diagrams and the corresponding singularities are displayed at each node (see Figs. 3(a) and 3(b), respectively); if ShowHistoryGraph -> SingOnly, only the singularities are shown.
@@
Finally, Substitutions allows the user to substitute some of the kinematic invariants (e.g., Substitutions -> {s12 -> 1}). Substitutions are applied before the polynomial equations are solved. V
@@
As mentioned in the discussion following (2.8), computing the maximal cut of a valid loop-by-loop Baikov representation may result in a degenerate (e.g., measure zero) integration contour. SOFIA[] implements heuristics to detect such situations and prints a warning. Whenever this happens, an alternative choice of LoopEdges might be needed; see examples in [repo]. In case of doubt, the warnings can be enabled by setting ShowPossiblyDegenerate -> True. V
Lines 13–21 in the code snippet above give options that control the solving strategy.
@@
Within [Uncaptioned image] , there are two available solvers: FastFubini (which implements the analytic procedure outlined in Sec. 2.2) and momentumPLD (which instead uses the numerical approach of Sec. 2.2). The former relies solely on Mathematica functions, while the latter requires the working Julia dependencies detailed in App. B. All the options PLDname relate specifically to momentumPLD (lines 13–18) are described under the same name as in [MathRepoTutorial]. Finally, the option FactorResult -> False can be used to skip the factorization into irreducible components of $\mathcal{S}(\vec{s})$ in (2.24), which, for large examples, can require both a long runtime and a large amount of memory.
We found that the native Mathematica solver outperforms the Julia solver on virtually all examples. Nevertheless, we keep the Julia solver available as an option to, e.g., cross-check the results obtained from the Fubini reduction in cases where it is practical. V
Lines 22–24 in the code snippet above set options for Effortless, which are relevant only when FindLetters -> True is enabled.
@@
The option AddLetters -> {{Even letters},{Odd letters}} allows the user to append additional even and odd letters to the list constructed by Effortless. The default value is the empty list {}. V
@@ 3.3.4 Graphical functions
The function FeynmanDraw opens a window that lets the user draw the desired Feynman diagram by clicking and dragging the pointer. It outputs a list of edges and nodes in the format {edges,nodes} with generic internal and external masses, which can be later adjusted by the user. By convention, the external momenta are all incoming and are assigned according to the ordering in nodes. The list {edges,nodes} completely specifies a Feynman diagram associated to (2.1), and is the main input of most functions in [Uncaptioned image] . V
@@
The Reflections -> True option leverages Symmetry[] (discussed below) to concurrently eliminate subtopologies related by reflection transformations (i.e., symmetries that do not involve any kinematic replacement rules). This option is particularly effective when most subtopologies are reflection-symmetric. For example, for the triple-box diagram in SOFIA_examples.nb [repo], this option is shown to improve the computation speed by approximately an order of magnitude. Note that behind the scenes, this is what the option Symmetries -> reflections is doing in the SOFIA[] command. V
@@
[…]
Figure 4: Diagrams related by the Symmetry[] function as described in the text. These diagrams were all generated by the FeynmanPlot[] function. The dashed lines denote massless particles and otherwise the mass label is displayed on the edges.
@@
In such cases, kinematic replacements can be subtle because the operations of specializing kinematics and computing singularities generally do not commute (a phenomenon analogous to that encountered when specializing kinematics and evaluating Feynman integrals; see [Mizera:2021icv, Dlapa:2023cvx, Fevola:2023kaw]). To address this, we introduce a small parameter $\delta$ and extract the leading non-vanishing coefficient for each component of the singular locus $\mathcal{S}$ as $\delta\to 0$. See App. C for an explicit demonstration.
Before moving on, we introduce the higher-level command SymmetryQuotient[L]. This command takes a list of diagrams formatted as L = {{edges1, nodes1}, {edges2, nodes2}, {edges3, nodes3}, ...} and returns a minimal set of representative diagrams, along with the corresponding replacement rules for kinematic scales needed to reproduce all the diagrams in the input list. In the main command SOFIA[], enabling the option Symmetries -> True internally applies SymmetryQuotient to the subtopologies generated by Subtopologies. V
@@ 4 Detailed examples
We now demonstrate the explicit use of the functions introduced above with concrete examples that are not too large to fit within the text. Collectively, these examples took less than 30 seconds to run on a conventional laptop. A list of 30 additional examples, including many multi-scale and multi-loop ones, can be found in the Mathematica notebook SOFIA_examples.nb [repo].
@@ 4.1 Symbol alphabet
In this example, we illustrate how [Uncaptioned image] , combined with the publicly available package Effortless [Effortlessxxx], can be used to determine a (super)set of logarithmic letters, both even and odd, for a given Feynman integral.
We consider the diagram in Fig. 5(a), which contributes to the $t$-channel planar Møller and Bhabha scattering amplitudes at $\mathcal{O}(e^{6})$ in QED. The massive electrons and/or positrons are represented by solid lines, while the massless photon exchanges are depicted by dashed lines. This diagram was computed explicitly using differential equations in [Duhr:2021fhk] and thus serves as an excellent crosscheck to the [Uncaptioned image] implementation.
@@
[…]
@@ 2 Sqrt[4 mm^3 - mm^2 s12 + 2 mm s12 s23 - s12 s23^2],
As we shall show momentarily, $\mathtt{r}$ is sufficient for the SOFIA command to produce a complete basis of odd letters. (Note that the signs chosen in the DoubleRoots option above are purely conventional and ensure that the comparison below is performed using the root conventions of [Duhr:2021fhk]).
At this stage, we can check the validity of SOFIAalphabet by comparing it against the alphabet knownAlphabet in [Duhr:2021fhk, Eq. (2.8)]. This is done by running:
@@
In line 1 above, the syntax reflects that the roots defined in [Duhr:2021fhk, Eq. (2.6)] were manually re-expressed, for comparison purposes, in terms of the roots $\mathtt{r[[i]]}$. The SOFIADecomposeAlphabet command then checks whether it is possible to decompose each element of knownAlphabet in terms of SOFIAalphabet and returns the decomposition if it is. For example, the most complicated letter in knownAlphabet decomposes according to:
$\hskip-8.5359pt\mathrm{d}\log\frac{r_{5}r_{10}\left(m^{2}{-}s_{12}\right){-}4m^{6}s_{12}{-}3m^{4}s_{12}^{2}{+}m^{2}s_{12}^{3}{+}3m^{2}s_{12}^{2}s_{23}{-}s_{12}^{3}s_{23}}{r_{5}r_{10}\left(m^{2}{-}s_{12}\right){+}4m^{6}s_{12}{+}3m^{4}s_{12}^{2}{-}m^{2}s_{12}^{3}{-}3m^{2}s_{12}^{2}s_{23}{+}s_{12}^{3}s_{23}}=\mathrm{d}\tilde{W}_{31}{-}2\mathrm{d}\tilde{W}_{32}\,.$ (4.1)
See the Mathematica notebook SOFIA_examples.nb [repo] for all the details of the complete decomposition, as well as how the roots $r_{i}=\mathtt{r[[i]]}$ in r and the letters $\tilde{W}_{j}$ in SOFIAalphabet are defined. In the same notebook, we show that the same procedure successfully reproduces other non-trivial alphabets, including those arising in the two-loop massless and one-mass pentagon functions [Chicherin:2017dob, Abreu:2020jxa, Chicherin:2021dyp] as well as the genuine two-loop massless six-particle letters [Abreu:2024fei, Henn:2025xrc].
@@ 4.2 Feynman geometries
An important early step in computing this amplitude is to determine the types of functions to which the contributing diagrams are expected to evaluate, i.e., the function space. This information is encoded in the underlying geometry of the associated master integrals. To reveal it, it suffices to carefully analyze the maximal cut of each master integral in four dimensions ($\varepsilon=0$), which is conveniently done in the Baikov representation thanks to (2.8). This approach is reliable because the underlying geometry is characterized by the solution of the homogeneous part of the differential equation satisfied by the uncut integrals, which coincides with the differential equation governing the maximal cut integral [Primo:2016ebd, Bosma:2017ens].
The maximal cut procedure always reduces Feynman integrals to periods over an algebraic variety. The geometry of this variety dictates the class of functions that appear after integration. For instance, a genus-0 (rational) curve gives rise to polylogarithms, a genus-1 (elliptic) curve to elliptic functions, and more intricate cases may involve higher-genus Riemann surfaces (see, e.g., [Marzucca:2023gto]) or even higher-dimensional Calabi–Yau manifolds (see Sec.,4.3), pointing to increasingly rich—and still not fully understood—mathematical structures. Strictly speaking, this way, one can only place an upper bound on the complexity of a given Feynman geometry; verifying that it is not degenerate generally requires further analysis. (To our knowledge, there is currently no automated algorithm to systematically determine the precise function space of a Feynman integral, e.g., whether it is polylogarithmic, elliptic, or beyond. Even less is known at the amplitude level, where full cancellations among non-polylogarithmic sectors may occur; see, e.g., [Bern:2024adl]).
@@
From (4.2), it is also easy to read off the leading Landau singularity (where all propagators are on-shell) of the diagram in Fig. 5(b). Indeed, these are given by the zeroes of the discriminant of $P_{4}(x_{7})$ in $x_{7}$, i.e.,
@@
One can verify that all the singularities in the $\varepsilon$-factorized differential equation from [Gorges:2023zgv] are included in the resulting list available in [repo].
@@ 4.3 Eikonal/Regge/heavy-mass physics
This example considers the eikonal/Regge scattering of two heavy masses—e.g., black holes or neutron stars, treated as point particles—via graviton exchange. In particular, we focus on the diagram shown in Fig. 5(c), which contributes to the process at $\mathcal{O}(G_{N}^{5})$. This example demonstrates two key points: first, how to take the eikonal limit of the Baikov representation; and second, how the latter turns out sufficient to uncover the underlying Calabi–Yau threefold geometry [Frellesvig:2023bbf, Frellesvig:2024rea].
In order to take the eikonal/Regge/heavy-mass expansion and keep the leading order term which captures the relevant classical physics [Brandhuber:2021eyq], we first parametrize the momenta in Fig. 5(c) such that
@@
The curve $\hat{P}_{4}(x_{12},x_{13},\hat{x}_{14})=0$ defines a Calabi–Yau threefold (as a quartic polynomial in $\mathbb{CP}^{3}$ [Hubsch:1992nu]), in agreement with the results of [Frellesvig:2023bbf, Frellesvig:2024rea].
@@ 4.4.1 Energy-energy correlators
As an example, we focus on the type of integrals arising in the computation of the four-point energy-energy correlator (EEC) in the collinear limit within $\mathcal{N}=4$ super Yang-Mills theory [Chicherin:2024ifn]. Physically, the EEC captures meaningful correlations of energy flux after the emission of four final-state particles from an initial energetic parton and is now an active topic of research in both theoretical and experimental high-energy physics, see, e.g.[Hofman:2008ar, Larkoski:2017jix].
For four-point EEC, the relevant integrals to compute originate from a phase-space integral over a squared $1\to 4$ amplitude [Chicherin:2024ifn, Eqs. (2-4)]. For our discussion, it is sufficient to quote the two families of integrals considered in [Chicherin:2024ifn]:
@@
We treat $z$, $\bar{z}$, $w$, $\bar{w}$ as independent variables. The parametrization of these families is overcomplete since, e.g., $a_{2}$ and $a_{3}$ in (4.10) cannot be simultaneously positive. This would introduce an overlapping pole $\frac{1}{x_{12}x_{23}}$ which is forbidden by a version of Steinmann relations. We thus find it simpler to perform the singularity analysis at the level of master integrals introduced in [Chicherin:2024ifn, App. B] (in the same way we typically do the analysis at the level of individual Feynman diagram topologies, as opposed to all integrals at a given loop order at the same time).
The complete analysis is given in [repo]. Below, we illustrate the procedure on a simple master integral
@@
{z,z¯, $\{z,\;\overline{z},\;1-z,\;1-\overline{z},\;z-\overline{z},\;1-z\bar{z},\;z+\bar{z}-z\bar{z},\;z+\bar{z}-2$$\bar{z}\}\,$. (4.14)
Looping this command over all the master integrals in [Chicherin:2024ifn, App. B] immediately reproduces the polynomial part of the alphabet quoted in that reference, namely
@@
returns a distinctly complicated letter (see SOFIA_examples.nb [repo] for details):
@@
It has an interesting geometric interpretation. As explained in [Chicherin:2024ifn], the solutions to the maximal cut conditions of the integrals in (4.16) (i.e., $s_{123}=s_{234}=x_{1234}=0$) define a singular cubic curve after projection onto the $(x_{2},x_{3})$-plane. Its roots a,b,ca\,,b\,,c, which give rise to the non-integer polynomial part of the alphabet listed in [Chicherin:2024ifn, Eq. (12)], satisfy